A garden table and a bench cost

$721

combined. The garden table costs

$71

more than the bench. What is the cost of the bench?


Answer by Boreal(15235)

bench=x
garden table=x+71
Their sum, x+x+71=2x+71=721
subtract 71 from both sides
2x=650
divide by 2 both sides
x=$325 bench ANSWER
x+71=$396 table
They add to $721